OpenStudy (anonymous):

Help me SOlve: If you were in a freely falling elevator and you dropped a pencil, it would hover in front of you. Is there a force of gravity acting on the pencil? Defend your answer.

OpenStudy (jfraser):

is there a force of gravity acting on you while you're in the elevator?

OpenStudy (unklerhaukus):

F = ma, if there is acceleration there is a force

OpenStudy (anonymous):

Force of gravity is always present. Since your in a free falling elevator, it means that your apparent weight is zero (which is Fn=normal force). proof: since a=g \(F_{net}=F_g-F_n\\F_n=F_g-F_{net}\\F_n=mg-ma\\F_n=mg-mg\\F_n=0 N\)

OpenStudy (anonymous):

*you're